OpenStudy (anonymous):

The basketball team scored 12 points the first game and 40 points the eighth game of the season. How many points will they earn with this pattern on the 13th game of the season?

OpenStudy (anonymous):

would you say this is an arithmetic sequence?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

I would say this is an arithmetic sequence. Since on the first game they scored 12 and in the 8th they scored 40. I still don't know exactly how you would know unless it's implied that it is such a problem: the equation is an=a1+(n-1)d where n = the interval in this case the nth game and d=distance or the difference in points

OpenStudy (anonymous):

think you can tackle it from here?

OpenStudy (anonymous):

n=13 ? d= 28/?

OpenStudy (anonymous):

internet problems sorry: and well first you'll have to solve for d, which I think you are on your track. and yes, it's 28/(n-1) but when solving for d, here we plugged in from n= 8; let me show you work unless you understand what I'm saying

OpenStudy (anonymous):

work please

OpenStudy (anonymous):

an=a1+(n-1)d a8=12 +(8-1)d 40=12+7d 28=7d 4=d

OpenStudy (anonymous):

all I did was plug in the known values that were given to us into the formula to find d, the unknown

OpenStudy (anonymous):

now that we have d, we can find a13 think that you could do that?

OpenStudy (anonymous):

whats a1 ? 13?

OpenStudy (anonymous):

and whats an ?

OpenStudy (anonymous):

an is like the value of the sequence a at the nth interval. and a1 is the value of the sequence at the first interval. a8 would be the value of the 8th interval. understand me?

OpenStudy (anonymous):

im getting 62 ? is that right ?

OpenStudy (anonymous):

that is close but no its not what you do?

OpenStudy (anonymous):

is it 64 ? or is that to high ?

OpenStudy (anonymous):

that's going up more, the answer is 60 but do you understand how to get it?

OpenStudy (anonymous):

yeah i know what i did wrong

OpenStudy (anonymous):

cool

OpenStudy (anonymous):

any other question you might have?